OpenStudy (anonymous):

[4_01]Which of the following is the slope between the points (0, -10) and (-4, 0)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

2/5 -2/5 -5/2 5/2

hartnn (hartnn):

can u solve now using formula for m ?

OpenStudy (anonymous):

ill try

hartnn (hartnn):

want a hint ?

OpenStudy (anonymous):

2/5 ?

hartnn (hartnn):

no.

OpenStudy (anonymous):

-2/5?

hartnn (hartnn):

m= (0-(-10))/(-4-0) = 10 / (-4) = -5/2 got this ?

OpenStudy (anonymous):

ohh i see why thank you.

hartnn (hartnn):

welcome ^_^